英語プレゼンテーション技法 1 単位 (選択)

Scientific Presentation Skills  カーペンターウォルター · 講師/ 機械工学科生産システム講座

【授業目的】The focus of this class is the development of communication and

presentation skills. 【授業概要】 Students will be provided with ample opportunities, both formal and

informal, to improve their ability to communicate. Thus, the basic approach to teaching this course is interactive in nature and students will be expected to actively participate during each class session. The goal is for students to gain experience, and therefore feel more confident and comfortable, in developing and presenting technical and scientific information.

【キーワード】 oral presentaion， question and answer 【先行科目】『実用技術英語』 (1.0) 【関連科目】『初級技術英語』 (0.5)，『中級技術英語』 (0.5)，『上級技術英語』 (0.5) 【履修要件】 [要件] 【履修上の注意】日亜 STC 学生は必ず受講すること．日亜 STC 学生でない学生が

この科目を受講する場合は自分の学科のコーディネータに連絡し受講許可を受けること．

【到達目標】 1. How to deliver a speech or a presentation in a professional environment

2. The basic steps necessary to develop a clearly understood presentation

3. How to prepare the content of a presentation or speech

4. Practical experience in public speaking

5. The fundamentals of communicating while using good ”body language” 【授業計画】

1. Fundamentals of Presentations(I): An easy yet effective ”three-step” approach to preparing either an oral or a written presentation.

2. Fundamentals of Presentations(II): An easy yet effective ”three-step” approach to preparing either an oral or a written presentation.

3. Fundamentals of Presentations(III): An easy yet effective ”three-step” approach to preparing either an oral or a written presentation.

4. Exercise with fundamentals of presentions

5. The importance of body language

6. Using effective body language to communicate with the target audience

7. Pronunciation practice (I): stress (accent) and enunciation

8. Pronunciation practice(II): stress (accent) and enunciation 9. How to incorporate visual aids and other media into the presentation; when

and how to use the chalkboard

10. How to incorporate visual aids and other media into the presentation; when to use a power point presentation (PPP)

11. How to evaluate the quality of a presentation as a means of improving one’s own skill in delivering a successful presentation

12. Presentation exercises

13. The importance of time: beginning on time and ending on time

14. The place and role of question and answer sessions: how to prepare and how to answer questions from the audience

15. Exercise for question and answer

16. Depending on the interests of the students enrolled in the course other possible topics to be covered could also include, for example: using the internet to conduct research, how to cite internet sources of information, what to avoid when using the internet for research purposes

【成績評価基準】 examination (40%) and presentation excercise (60%)

【学習教育目標との関連】 [JABEE 関連]

【教科書】 Presenting Science: A Technical English Course for Engineering Majors.

(Second Edition). 2008. Macmillan LanguageHouse. 【参考書】 [参考資料] 【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=215664

【対象学生】 Nichia STC students

【連絡先】 ⇒ カーペンター (088-656-7643, walter@cicee.tokushima-u.ac.jp) ���� ⇒ 橋爪 (E 棟 3 階南 D-2, 088-656-7473, tume@ee.tokushima-u.ac.jp) ����

Scientific Presentation Skills 1 unit (selection)

Walter Carpenter · Associate Professor / Production Systems Engineering, Department of Mechanical Engineering

Target⟩ The focus of this class is the development of communication and

presentation skills. Outline⟩ Students will be provided with ample opportunities, both formal and

informal, to improve their ability to communicate. Thus, the basic approach to teaching this course is interactive in nature and students will be expected to actively participate during each class session. The goal is for students to gain experience, and therefore feel more confident and comfortable, in developing and presenting technical and scientific information.

�������⟩ oral presentaion, question and answer Fundamental Lecture⟩ “Practical Technical English”(1.0) Relational Lecture⟩ “Basic Technical English”(0.5), “Technical English”(0.5),

“Advanced Technical English”(0.5) Requirement⟩ [要件] Notice⟩ A limited number of students will be permitted to take this class.This

course is a compulsory one for Nichia STC students. If you are not a Nichia STC student and want to attend this class, please contact a coordinator in your department for getting permission for your attendance.

Goal⟩

1. How to deliver a speech or a presentation in a professional environment

2. The basic steps necessary to develop a clearly understood presentation

3. How to prepare the content of a presentation or speech

4. Practical experience in public speaking

5. The fundamentals of communicating while using good ”body language”

Schedule⟩

1. Fundamentals of Presentations(I): An easy yet effective ”three-step” approach to preparing either an oral or a written presentation.

2. Fundamentals of Presentations(II): An easy yet effective ”three-step” approach to preparing either an oral or a written presentation.

3. Fundamentals of Presentations(III): An easy yet effective ”three-step” approach to preparing either an oral or a written presentation.

4. Exercise with fundamentals of presentions

5. The importance of body language

6. Using effective body language to communicate with the target audience 7. Pronunciation practice (I): stress (accent) and enunciation

8. Pronunciation practice(II): stress (accent) and enunciation

9. How to incorporate visual aids and other media into the presentation; when and how to use the chalkboard

10. How to incorporate visual aids and other media into the presentation; when to use a power point presentation (PPP)

11. How to evaluate the quality of a presentation as a means of improving one’s own skill in delivering a successful presentation

12. Presentation exercises

13. The importance of time: beginning on time and ending on time

14. The place and role of question and answer sessions: how to prepare and how to answer questions from the audience

15. Exercise for question and answer

16. Depending on the interests of the students enrolled in the course other possible topics to be covered could also include, for example: using the internet to conduct research, how to cite internet sources of information, what to avoid when using the internet for research purposes

Evaluation Criteria⟩ examination (40%) and presentation excercise (60%)

Relation to Goal⟩ [JABEE 関連]

Textbook⟩ Presenting Science: A Technical English Course for Engineering

Majors. (Second Edition). 2008. Macmillan LanguageHouse. Reference⟩ [参考資料] Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=215664

Student⟩ Nichia STC students

Contact⟩ ⇒ Carpenter (+81-88-656-7643, walter@cicee.tokushima-u.ac.jp) ���� ⇒ Hashizume (E 棟 3 階南 D-2, +81-88-656-7473, tume@ee.tokushima-u.ac.j

p) ����